OpenStudy (anonymous):

The perimeter of a rectangle is the sum of all of its sides. Write a mathematical expression for the perimeter of a rectangle with a length of 4 and a width of w. Simplify the expression if possible.

OpenStudy (anonymous):

4+4+w+w = 8+2w 8+2w= Perimeter of rectangle